OpenStudy (anonymous):

y=(x+3)^-1 in to standard form.. how do u do that\/?

OpenStudy (anonymous):

you know that \[x ^{-1}\] = 1/x

OpenStudy (anonymous):

im unsure of what standard form means here..do you have a basic example of and equation in standard form? that equation above is y=1/(x+3) im not sure if thats the answer

OpenStudy (anonymous):

do you mean y=mx+b?

OpenStudy (anonymous):

i wrote it lik y+(x+3)(x+3)-1

OpenStudy (anonymous):

how did yo uget that?